# GRPO-LEAD: A Difficulty-Aware Reinforcement Learning Approach for Concise Mathematical Reasoning in Language Models

### v1

@@ Abstract
Recent advances in R1-like reasoning models leveraging Group Relative Policy Optimization (GRPO) have significantly improved the performance of language models on mathematical reasoning tasks. However, current GRPO implementations encounter critical challenges, including reward sparsity due to binary accuracy metrics, limited incentives for conciseness, and insufficient focus on complex reasoning tasks. To address these issues, we propose GRPO-LEAD, a suite of novel enhancements tailored for mathematical reasoning. Specifically, GRPO-LEAD introduces (1) a length-dependent accuracy reward to encourage concise and precise solutions, (2) an explicit penalty mechanism for incorrect answers to sharpen decision boundaries, and (3) a difficulty-aware advantage reweighting strategy that amplifies learning signals for challenging problems. Furthermore, we systematically examine the impact of model scale and supervised fine-tuning (SFT) strategies, demonstrating that larger-scale base models and carefully curated datasets significantly enhance reinforcement learning effectiveness. Extensive empirical evaluations and ablation studies confirm that GRPO-LEAD substantially mitigates previous shortcomings, resulting in language models that produce more concise, accurate, and robust reasoning across diverse mathematical tasks. Our source code, generated dataset, and model are available at https://github.com/aeroplanepaper/GRPO-LEAD.
@@ 1 Introduction
Figure 1: Validation Pass@1 over training steps for three configurations: GRPO, GRPO with length reward, and GRPO with length reward plus advantage reweighting. The validation consists of 27 challenging problems from AIMO2 Frieder et al. (2024), CMU-MATH-AIMO’s validation Sun (2024), and AIME24, where Deepseek-14B struggles to solve.
Recently, R1-like reasoning models have attracted significant attention due to their impressive performance in solving challenging mathematical reasoning tasks through extensive chains of thoughtLuo et al. (2025); Wen et al. (2025). According to the technical report introducing R1Guo et al. (2025), reinforcement learning (RL) fine-tuning plays a pivotal role in enabling this reasoning capability. In particular, Group Relative Policy Optimization (GRPO)Shao et al. (2024), a novel RL approach for language models, has emerged as a promising alternative to traditional methods such as PPOSchulman et al. (2017) and DPORafailov et al. (2023), primarily due to its efficiency and intrinsic compatibility with language model training.
However, existing GRPO implementations still encounter substantial limitations. One key issue is the inherent reward sparsity arising from binary and rule-based accuracy metrics, which significantly hampers effective model training. Specifically, when all generated responses to a given problem are either uniformly correct or incorrect, the resulting uniform reward signal provides minimal differentiation, leading to weak learning gradients and consequently slower model convergence. To elaborate, if all outputs in a question group are correct, each receives identical positive feedback, diluting the informative gradient needed for meaningful policy improvements. Conversely, uniformly incorrect responses yield no useful information to guide policy refinement.
@@
Motivated by these limitations, this work introduces GRPO-LEAD, a suite of targeted modifications explicitly designed to enhance GRPO’s effectiveness for mathematical reasoning tasks. Our key contributions include:
  1. 1.
    Length-dependent Accuracy Reward: Introduces a dynamic reward shaping mechanism that promotes brevity among correct responses using standardized length-based penalties, reducing verbosity without sacrificing accuracy.
  2. 2.
    Explicit Penalty for Incorrect Solutions: Implements a negative reward for incorrect outputs to enforce a sharper decision boundary, mitigating overconfidence and boosting precision of the model.
  3. 3.
    Difficulty-aware Advantage Reweighting: Applies a logistic weighting function to advantage estimates based on empirical correctness rates, ensuring stronger updates for harder problems and balanced generalization.
  4. 4.
    Impact of Model Scale and Data Quality on Reinforcement Learning Effectiveness: Demonstrates that larger base models and high-quality, curriculum-structured fine-tuning data significantly improve RL convergence and output quality; also introduces targeted reward interventions to mitigate repetitive formatting artifacts.
Empirical evaluations and comprehensive ablation studies confirm that our method effectively addresses previous GRPO shortcomings, leading to more concise, accurate, and efficiently trainable models capable of robust performance across mathematical reasoning tasks.
@@ 2.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO), a recently proposed algorithm designed specifically for fine-tuning language models with group-level normalization of rewardsGuo et al. (2025). GRPO modifies the standard policy gradient objective by introducing a relative advantage within sets of responses corresponding to the same question, stabilizing updates, and promoting consistent learning signals. Formally, GRPO defines the objective as:
$\mathcal{L}_{\text{GRPO}}(\theta)=-\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigg{[}$ $\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}\hat{A}_{i,t}$
$\quad-\beta D_{\text{KL}}[\pi_{\theta}\|\pi_{\theta_{\text{old}}}]\Bigg{]},$ (1)
where $G$ is the number of groups, $\hat{A}_{i,t}$ is the normalized advantage within the group, and $\beta$ controls the KL divergence penalty term enforcing policy stability.
@@ 2.2 Length Reward
A prevalent issue in reinforcement learning-based fine-tuning of language models is the uncontrolled increase in response length driven by reward signals, commonly known as reward hacking. This phenomenon leads to unnecessarily verbose responses, which, although technically correct, often lack conciseness and hinder interpretability. Furthermore, such verbosity fails to reflect efficient reasoning, limiting model utility in practical scenarios. Existing efforts to mitigate this problem typically involve incentivizing shorter answers to encourage more succinct reasoning processes. For example, Kimi proposed an individual min-max normalized length reward based on the lengths of generated responses Team et al. (2025). Yeo et al. introduced a cosine length reward function with fixed maximum and minimum thresholds to manage response lengths Yeo et al. (2025). Aggarwal et al. utilized a target "golden length" to directly reward or penalize responses based on their deviation from an ideal length Aggarwal and Welleck (2025).
@@ 3 Method
To systematically address the limitations identified in existing implementations of Group Relative Policy Optimization (GRPO), we propose a suite of novel modifications collectively termed GRPO-LEAD (GRPO with Length-dependent rewards, Explicit penalties, and Advantage reweighting for Difficulty). Our proposed method enhances the original GRPO framework by introducing three core innovations: 1) a length-dependent accuracy reward to foster concise solutions, 2) an explicit penalty mechanism to mitigate low precision rate caused by length reward, and 3) an advantage reweighting strategy tailored to compensate for limited rollout scenarios. Additionally, we rigorously examine how base model scale and supervised fine-tuning (SFT) impact the effectiveness of reinforcement learning (RL) fine-tuning.
@@ 3.1 Length-Dependent Accuracy Reward
3.2 Explicit Penalty for Incorrect Answers
Existing methods often focus on maximizing pass@1 within a restricted context length. However, we observe that prioritizing pass@1 can reduce overall accuracy. Our experiments indicate that this decline is not directly caused by incorporating a length reward but rather by relying on a binary accuracy reward.
When length-based regularization is applied, it may shorten responses and limit the model’s ability to “re-think” or revise incorrect answers through longer reasoning. Still, disabling the length reward does not fully address the problem. Although pass@1 improves for both training and evaluation, the solve-all metric—measured by how many questions are answered correctly in all sampled responses—tends to drop. Adding the length reward simply accelerates this trend.
We attribute this outcome to the binary reward signal. Under a limited generation length, a model is incentivized to complete its response within the available budget. While such an approach may be less rigorous and more prone to errors, it can still yield a non-zero reward, since providing a guess is better than providing none at all. This creates a shortcut, where ambiguous or partially complete answers inflate pass@1 yet lower overall precision.
To mitigate this, we propose a revised reward structure that explicitly penalizes incorrect responses, thereby reinforcing a sharper decision boundary between correct and incorrect outputs. The revised reward function is defined as:
@@ 3.2 Explicit Penalty for Incorrect Answers
where $z$ denotes the standardized length deviation, and $\alpha>0$ controls the strength of length penalization for correct responses, as previously defined.
Under this formulation, the expected reward for a response with correctness probability $P(\text{correct})$ is:
@@
To gain intuition about the behavior of this reward function, we consider a simplified case where the length penalty is neutralized, i.e., $\exp(-\alpha z)\approx 1$(In practice the mean reward of correct answer is approcximately 1). Under this assumption:
@@
This approximation highlights a key property of the reward: the expected value becomes positive only when $P(\text{correct})>0.5$.
This formulation introduces a principled deterrent against random guessing and encourages the model to internalize a more robust decision threshold. Empirically, this approach improves both pass@1 and overall precision.
@@ 3.3 Advantage Reweighting for Difficulty-Aware Training
While the length-dependent accuracy reward and explicit penalty formulation described earlier significantly mitigate verbosity and enhance precision, a subtle yet critical challenge remains. In the original GRPO algorithm, the rewards are uniformly applied across questions regardless of intrinsic difficulty. This uniformity implicitly biases the model to excessively optimize performance on simpler tasks—where obtaining correct and concise responses is inherently easier—while inadvertently neglecting harder questions that require deeper reasoning and longer deliberation. Consequently, the model may degrade on complex problems due to an insufficient incentive structure.
To address this imbalance, we introduce a difficulty-aware advantage reweighting strategy designed to dynamically adjust the magnitude of policy updates based on problem difficulty. Intuitively, we aim to amplify learning signals for harder problems, thereby encouraging sustained exploration and deeper reasoning strategies rather than superficial optimization on trivial tasks.
Formally, given a prompt set grouped by question $q$ with associated responses ${o_{i}}$, we first define the group’s empirical correctness ratio:
$\rho_{q}=\frac{\text{number of correct responses for }q}{\text{total number ofresponses for }q}.$ (5)
This correctness ratio serves as a proxy for problem difficulty, with lower values indicating greater difficulty.
@@
To apply this weighting to the advantage calculation, consider the normalized advantage estimate for response $o_{i}$:
$\tilde{A}i=\frac{R(o_{i}|q)-\mu_{q}}{\sigma_{q}+\epsilon},$ (7)
where $\mu_{q}$ and $\sigma_{q}$ represent the mean and standard deviation of rewards within the prompt group $q$. We then define the difficulty-aware advantage as:
@@
3.4 Impact of Model Scale and Data Quality on Reinforcement Learning Effectiveness
While the preceding reward strategies effectively address verbosity and improve precision across varying difficulties, the base model’s scale and the quality of the underlying dataset significantly influence the effectiveness of RL process. Specifically, we analyze how initial model capability and carefully curated training data jointly affect RL fine-tuning outcomes.
Firstly, our experiments reveal a clear dependence between model scale and RL improvements. While RL fine-tuning notably enhances the Deepseek Distilled 7B model’s performance on relatively straightforward questions, substantial gains on complex reasoning tasks remain elusive. The 7B model often prematurely converges to incorrect reasoning paths and frequently overlooks critical edge cases. Conversely, we observe that the Deepseek Distilled 14B model inherently outperforms even the RL-enhanced 7B variant, particularly on questions requiring exhaustive reasoning or enumeration of multiple scenarios. These findings suggest intrinsic limitations in smaller models’ capacities for advanced reasoning, underscoring the importance of larger-scale base models for RL fine-tuning to effectively tackle complex problems.
To further investigate methods for enhancing model capability, we generated a targeted dataset comprising 13k math reasoning problems sourced from the dataset provided by DeepScalerLuo et al. (2025), including historical AMC, AIME, and OmniMath exams, accompanied by solutions generated using QwQ32BTeam (2025), a state-of-the-art small scale reasoning model. After supervised fine-tuning (SFT) with this specialized dataset, we applied our proposed RL strategies. Despite initial signs of overfitting in the SFT stage, we observed that subsequent RL fine-tuning rapidly alleviated these issues, demonstrating faster convergence and significant improvements in both pass@1 accuracy and overall precision relative to RL fine-tuning applied directly to the original model.
Our results additionally highlight the critical role of data quality and curriculum strategies in sustained RL improvement. By first applying RL on a subset of approximately 7k challenging problems (difficulty rating $\geq 4$ from the DeepScaler dataset), we obtained a robust initial policy checkpoint. Subsequently, we refined this policy using a curriculum consisting of the most challenging problems identified from the first-stage correctness distribution (correctness rate below 50%), supplemented by high-difficulty examples from the second stagen dataset of Light-R1 Wen et al. (2025). Empirical evaluations demonstrate that this two-stage curriculum significantly enhances the model’s ability to continuously improve on complex tasks.
Finally, we addressed a persistent formatting issue—repetitive n-gram patterns likely arising from the absence of clear end-of-sequence (EOS) signals during the initial SFT stage. By temporarily removing length-dependent rewards and introducing an explicit negative reward ($-1.5$) for repeated n-grams, we achieved further improvements in precision and pass@1 metrics. This intervention demonstrates the effectiveness of targeted reward modifications in mitigating specific output-formatting anomalies.
In summary, our experiments affirm the substantial influence of initial model capacity, dataset quality, and targeted reward engineering on RL-based fine-tuning outcomes. These findings collectively inform a systematic approach for enhancing language models’ capability to produce concise, accurate, and well-structured responses across tasks of varying complexity.
@@ 4.1 Datasets and Filtering
Our primary training data is sourced from the DeepScaler dataset Luo et al. (2025). We begin by filtering out problems with difficulty ratings below 2.5, resulting in approximately 9,000 questions for fine-tuning.
For stages 2 and 3 of our 14B model experiments, we further refine the dataset by selecting problems where the model’s stage-1 rollout accuracy is no greater than 75%, yielding around 2,283 questions. Additionally, we incorporate challenging problems with numeric answers from the stage-2 dataset of Light-R1 Wen et al. (2025).
In total, the dataset for stages 2 and 3 comprises 3,524 questions. This adaptive filtering strategy ensures a focused emphasis on harder problems, aiming to improve the model’s performance on more complex tasks.

### v2

@@ Abstract
Group Relative Policy Optimization (GRPO), which is widely adopted by R1-like reasoning models, has advanced mathematical reasoning. Nevertheless, GRPO faces challenges in reward sparsity, verbosity, and inadequate focus on problem difficulty. We propose GRPO-LEAD, enhancing GRPO with: (1) length-regularized rewards to encourage conciseness while maintaining accuracy; (2) explicit penalties for incorrect solutions to improve model precision; and (3) difficulty-aware advantage reweighting for robust generalization on challenging problems. Comprehensive evaluations demonstrate that GRPO-LEAD significantly improves reasoning accuracy, conciseness, and efficiency. Our approach achieves state-of-the-art performance for 14B-scale models, underscoring the synergy of our methods with appropriate model scale and high-quality data. Our source code, generated dataset, and models are available at https://github.com/aeroplanepaper/GRPO-LEAD.
GRPO-LEAD: A Difficulty-Aware Reinforcement Learning Approach for Concise Mathematical Reasoning in Language Models
Jixiao Zhang [^1] Equal contribution. Johns Hopkins University jzhan432@jh.edu          Chunsheng Zuo [^1] Equal contribution. Johns Hopkins University czuo3@jh.edu
@@ 1 Introduction
Figure 1: The GRPO-LEAD framework assigns length-regularized positive rewards to correct answers and explicit penalties to incorrect ones. A difficulty-based weight $w$ used for advantage reweighting is determined from the empirical correctness of responses for each question. This weight then scales the advantages derived from each question, prioritizing harder questions over easier ones during the policy update to foster robust reasoning.
Recently, R1-like reasoning models have attracted significant attention due to their impressive performance in solving challenging mathematical reasoning tasks through extensive chains of thought Luo et al. (2025b); Wen et al. (2025). According to the technical report introducing R1 Guo et al. (2025), reinforcement learning (RL) fine-tuning plays a pivotal role in enabling this reasoning capability. In particular, Group Relative Policy Optimization (GRPO) Shao et al. (2024), a novel RL approach for language models, has emerged as a promising alternative to traditional methods such as PPO Schulman et al. (2017) and DPO Rafailov et al. (2023), primarily due to its efficiency and intrinsic compatibility with language model training. Researchers across various domains have successfully employed GRPO Li et al. (2025); Liu et al. (2025a); Luo et al. (2025a); Dai et al. (2025), achieving impressive outcomes.
Despite its strengths, existing GRPO implementations encounter significant limitations. A primary issue is reward sparsity stemming from binary, rule-based accuracy metrics; when responses within problem groups exhibit uniform correctness or incorrectness, the resulting uniform reward signals offer minimal differentiation, weakening learning gradients and hampering convergence. Moreover, such uniform signals inadequately promote concise reasoning, leading to unnecessarily verbose outputs and inefficiencies during training and inference. Additionally, the current reward formulation lacks explicit penalties for incorrect answers Hu et al. (2025a); Luo et al. (2025b); Chu et al. (2025), inadvertently encouraging models to guess rather than engage in rigorous reasoning, thereby compromising precision. Furthermore, rewards are applied uniformly across problems regardless of their intrinsic difficulty, causing models to excessively optimize simpler tasks while neglecting more challenging problems that require deeper reasoning.
@@
Motivated by these limitations, this work introduces GRPO-LEAD, a suite of targeted modifications explicitly designed to enhance GRPO’s effectiveness for mathematical reasoning tasks. The overall framework is illustrated in Figure 1. Our key contributions include:
  • We introduce a length-regularized reward with an explicit penalty for incorrect solutions to encourage solution conciseness while maintaining accuracy.
  • We apply difficulty-aware advantage reweighting to focus learning on more challenging problems, fostering robust generalization.
  • Our comprehensive evaluations demonstrate GRPO-LEAD significantly improves reasoning accuracy and conciseness, achieving state-of-the-art performance in mathematical reasoning for 14B-scale models.
@@ 2.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) is a recently proposed algorithm designed specifically for fine-tuning language models with group-level normalization of rewards Guo et al. (2025). GRPO modifies the standard policy gradient objective by introducing relative advantages within sets of responses corresponding to the same query, stabilizing updates and promoting consistent learning signals. Formally, GRPO defines the objective as:
$\mathcal{L}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Big[\min\big(r_{i,t}(\theta)\hat{A}_{i,t},\,$ (1)
$\text{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\big)\Big]$
where the importance sampling ratio is given by
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$ (2)
Here, $G$ denotes the number of groups (e.g., different queries), $\hat{A}_{i,t}$ is the normalized advantage within group $i$, and $\epsilon$ defines the clipping range for conservative updates.
@@ 2.2 Length Reward
A prevalent issue in reinforcement learning-based fine-tuning of language models is reward hacking Everitt et al. (2017); Gao et al. (2023); Weng (2024). In GRPO, when the model is trained with a large fixed budget, it can exploit this budget by producing an excessive number of extra reasoning and verification steps to ensure the correctness of the answer and therefore reach a higher reward. This phenomenon leads to unnecessarily verbose responses that lack conciseness and hinder interpretability, resulting in inefficiency in reasoning and reducing the model’s practicality.
Existing efforts to mitigate this problem typically involve incentivizing shorter answers to encourage more succinct reasoning processes. For example, Kimi proposed an individual min-max normalized length reward based on the lengths of generated responses Team et al. (2025). Yeo et al. introduced a cosine length reward function with fixed maximum and minimum thresholds to manage response lengths Yeo et al. (2025). Aggarwal et al. utilized a target "golden length" to directly reward or penalize responses based on their deviation from an ideal length Aggarwal and Welleck (2025).
@@ 3 Method
To systematically address the limitations identified in existing implementations of Group Relative Policy Optimization (GRPO), we propose a suite of novel modifications collectively termed GRPO-LEAD (GRPO with Length-dependent rewards, Explicit penalties, and Advantage reweighting for Difficulty). Our proposed method enhances the original GRPO framework by introducing three core innovations: 1) a length-dependent accuracy reward to foster concise solutions, 2) an explicit penalty mechanism to mitigate low precision rate caused by length reward, and 3) a difficulty-aware advantage reweighting strategy that amplifies learning signals for challenging problems. Additionally, we examine how base model scale and supervised fine-tuning (SFT) impact the effectiveness of reinforcement learning (RL) fine-tuning.
@@ 3.1 Length-Dependent Accuracy Reward
3.2 Explicit Penalty for Incorrect Answers to Enhance True Accuracy
Existing methods often prioritize maximizing pass@1—the success rate on the first attempt—typically within restricted response lengths. However, this focus can inadvertently degrade overall model accuracy. The fundamental issue appears to stem from the use of a binary accuracy reward, rather than length-based regularization: under pressure to generate responses within a limited length, a model is encouraged to provide an answer, even if it’s a guess, rather than no answer at all. Such guesses can achieve a non-zero reward and inflate pass@1, but they do so at the cost of overall precision by rewarding less rigorous reasoning.
To counteract this tendency and foster a more robust distinction between correct and incorrect outputs, we introduce a revised reward structure that explicitly penalizes incorrect responses. This new reward function is defined as:
@@ 3.2 Explicit Penalty for Incorrect Answers to Enhance True Accuracy
where $o$ is the output, $q$ is the question, $z$ represents the standardized length deviation of a correct response, and $\alpha>0$ is a hyperparameter controlling the strength of the length penalization for correct answers, consistent with prior definitions.
The expected reward for a response, given its probability of correctness $P(\text{correct})$, under this formulation is:
@@
To intuitively grasp the impact of this reward function, let us consider a simplified scenario where the length penalty for correct answers is negligible (i.e., $\exp(-\alpha z)\approx 1$). In practice, the average reward for correct answers often normalizes close to this value. Under this assumption, the expected reward simplifies to:
@@
This approximation reveals a crucial characteristic: the expected reward becomes positive only when $P(\text{correct})>0.5$. This threshold acts as a principled deterrent against speculative guessing, compelling the model to internalize a more stringent decision boundary for correctness. Our empirical results confirm that this approach significantly improves both pass@1 and overall precision, encouraging the model to favor accuracy over mere completion.
@@ 3.3 Advantage Reweighting for Difficulty-Aware Training
While length reward and advantage reweighting can enhance precision and mitigate verbosity, uniformly applying rewards across all questions, irrespective of their intrinsic difficulty, may implicitly bias the model. It might learn to excessively optimize performance on simpler tasks-—where correct and concise responses are more readily achieved-—while neglecting more complex questions that demand deeper reasoning. Consequently, the performance on challenging problems can degrade.
Therefore, we introduce a difficulty-aware advantage reweighting strategy, which dynamically adjusts the magnitude of policy updates based on an estimate of problem difficulty. The intuition is to amplify learning signals for harder tasks, re-anchoring the model towards harder tasks.
Formally, we first quantify problem difficulty. For a given question $q$ and its associated set of sampled responses $\{o_{i}\}$, we define the group’s empirical correctness ratio as:
[…]
This ratio, $\rho_{q}$, serves as an inverse proxy for problem difficulty: a lower $\rho_{q}$ suggests a harder question.
@@
To apply this reweighting, we first consider the normalized advantage estimate for a response $o_{i}$ to question $q$:
$\tilde{A}_{i}=\frac{R(o_{i}|q)-\mu_{q}}{\sigma_{q}+\epsilon},$ (8)
where $\mu_{q}$ and $\sigma_{q}$ are the mean and standard deviation of rewards $R(o_{i}|q)$ for responses to question $q$, and $\epsilon$ is a small constant for numerical stability. We then define the difficulty-aware advantage, $A_{i}^{\prime}$, as:
@@
3.4 Impact of Data Quality on Reinforcement Learning Effectiveness
To further enhance model capabilities, we first performed supervised fine-tuning (SFT) on a specialized dataset of 13k math reasoning problems sourced from DeepScaler Luo et al. (2025b) (including historical AMC, AIME, and OmniMath problems) with solutions generated by QwQ32B Team (2025). Although this SFT model initially showed signs of overfitting, subsequent application of our proposed RL strategies rapidly mitigated these issues. This SFT+RL approach yielded faster convergence and significantly improved pass@1 accuracy and overall precision compared to applying RL directly to the original base model.
Our findings also highlight the critical role of data quality and curriculum strategies in RL. We established a robust initial policy by applying RL to a subset of challenging problems from the DeepScaler dataset. This policy was then further refined using a curriculum composed of the most challenging problems identified from this first RL stage and supplemented by high-difficulty examples from the Light-R1 dataset Wen et al. (2025). This two-stage curriculum markedly enhanced the model’s ability to continuously improve on complex tasks.
Finally, we addressed a persistent formatting issue of repetitive n-gram patterns, likely stemming from an absence of clear end-of-sequence (EOS) signals during SFT. By temporarily removing length-dependent rewards and introducing an explicit negative reward (1.5-1.5) for such repeated n-grams, we achieved further improvements in precision and pass@1 metrics. This intervention demonstrates the effectiveness of targeted reward modifications for mitigating specific output anomalies.
In summary, our experiments affirm that initial model capacity, curated data curricula for RL, and targeted reward engineering are pivotal for optimizing fine-tuning outcomes. These elements collectively inform a systematic approach for enhancing language models’ ability to produce concise, accurate, and well-structured responses across tasks of varying complexity.
@@ 4.1 Datasets and Filtering
Our primary training data is sourced from the DeepScaler dataset Luo et al. (2025b). We filter out problems with difficulty ratings below 2.5, resulting in approximately 9,000 questions for fine-tuning.
For stage 2 of our 14B model experiments, we further refine the dataset by selecting problems where the model’s stage-1 rollout accuracy is no greater than 75%, yielding around 2,283 questions. Additionally, we incorporate challenging problems with numeric answers from the stage-2 dataset of Light-R1 Wen et al. (2025), resulting in 3,524 questions in total.